OpenStudy (anonymous):

Estimate graphically the local maximum and local minimum of f(x) = 4x2 + 3x + 2 Local maximum: (-0.4,1.44); local minimum: (0,-0.37) Local maximum: (-0.4,1.44); no local minimum No local maximum; local minimum: (-0.4,1.44) No local maximum; local minimum: (0,-0.37)

OpenStudy (anonymous):

\[ f(x) = 4x^2 + 3x + 2 \] a parabola that opens up, so no max but a min at the vertex

OpenStudy (precal):

use x=-b/(2a) to figure out your vertex value

OpenStudy (anonymous):

when i did ut i got c

OpenStudy (anonymous):

vertex at \(x=-\frac{b}{2a}=-\frac{3}{8}\)

OpenStudy (anonymous):

The functionfunction A set of ordered pairs where each first element is paired with one and only one second element and no element in either pair is without a partner. is neither even nor odd

OpenStudy (anonymous):

!!so D is correct

OpenStudy (anonymous):

idrk this oneee but take a guess :)

OpenStudy (anonymous):

no i think it might be C it is usually C

OpenStudy (anonymous):

@satellite73 okkkk

OpenStudy (anonymous):

A plot of\[4 x^2+3 x+2 \] where x ranges from -0.6 through -.02 is attached. The calculus indicates that (-0.375,1.4375) is the local minimum.

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

aww np